# High-Strength Amorphous Silicon Carbide for Nanomechanics

### v1

@@ Acknowledgements.
- Barboni et al. [2018] L. Barboni, G. Gillich, C. Chioncel, C. Hamat, and I. Mituletu, A method to precise determine the young’s modulus from dynamic measurements, in IOP Conference Series: Materials Science and Engineering, Vol. 416 (IOP Publishing, 2018) p. 012063.
- Chirikov et al. [2020] V. A. Chirikov, D. M. Dimitrov, and Y. S. Boyadjiev, Determination of the dynamic young’s modulus and poisson’s ratio based on higher frequencies of beam transverse vibration, Procedia Manufacturing 46, 87 (2020).
- Chen [2000] S. Chen, Resonant frequency method for the measurement and uncertainty analysis of acoustic and elastic properties, Ultrasonics 38, 206 (2000).
- Villanueva and Schmid [2014a] L. G. Villanueva and S. Schmid, Evidence of surface loss as ubiquitous limiting damping mechanism in sin micro-and nanomechanical resonators, Physical review letters 113, 227201 (2014a).
- Wang et al. [2017] S. Wang, Z. Shan, and H. Huang, The mechanical properties of nanowires, Advanced Science 4, 1600332 (2017).
- Banerjee et al. [2018] A. Banerjee, D. Bernoulli, H. Zhang, M.-F. Yuen, J. Liu, J. Dong, F. Ding, J. Lu, M. Dao, W. Zhang, et al., Ultralarge elastic deformation of nanoscale diamond, Science 360, 300 (2018).
@@
- Shafikov et al. [2021] A. Shafikov, B. Schurink, R. W. van de Kruijs, J. Benschop, W. Van den Beld, Z. S. Houweling, and F. Bijkerk, Strengthening ultrathin si3n4 membranes by compressive surface stress, Sensors and Actuators A: Physical 317, 112456 (2021).
@@
- Demetriou et al. [2011] M. D. Demetriou, M. E. Launey, G. Garrett, J. P. Schramm, D. C. Hofmann, W. L. Johnson, and R. O. Ritchie, A damage-tolerant glass, Nature materials 10, 123 (2011).
- Zhang et al. [2022] S. Zhang, Z. Li, K. Luo, J. He, Y. Gao, A. V. Soldatov, V. Benavides, K. Shi, A. Nie, B. Zhang, et al., Discovery of carbon-based strongest and hardest amorphous material, National Science Review 9, nwab140 (2022).
- Namazu [2023] T. Namazu, Mechanical property measurement of micro/nanoscale materials for mems: A review, IEEJ Transactions on Electrical and Electronic Engineering 18, 308 (2023).
- Grunenberg [2001] J. Grunenberg, Intrinsic bond strengths of multiple c- c, si- si, and c- si bonds, Angewandte Chemie International Edition 40, 4027 (2001).
- Villanueva and Schmid [2014b] L. G. Villanueva and S. Schmid, Evidence of surface loss as ubiquitous limiting damping mechanism in sin micro-and nanomechanical resonators, Physical review letters 113, 227201 (2014b).
@@
- Fedorov et al. [2019] S. A. Fedorov, N. J. Engelsen, A. H. Ghadimi, M. J. Bereyhi, R. Schilling, D. J. Wilson, and T. J. Kippenberg, Generalized dissipation dilution in strained mechanical resonators, Physical Review B 99, 054107 (2019).
- Guo et al. [2019] J. Guo, R. Norte, and S. Gröblacher, Feedback cooling of a room temperature mechanical oscillator close to its motional ground state, Physical review letters 123, 223602 (2019).
- González and Saulson [1994] G. I. González and P. R. Saulson, Brownian motion of a mass suspended by an anelastic wire, The Journal of the Acoustical Society of America 96, 207 (1994).
- Tsaturyan et al. [2017] Y. Tsaturyan, A. Barg, E. S. Polzik, and A. Schliesser, Ultracoherent nanomechanical resonators via soft clamping and dissipation dilution, Nature nanotechnology 12, 776 (2017).
@@
- Li et al. [2023] Z. Li, M. Xu, R. A. Norte, A. M. Aragón, F. Van Keulen, F. Alijani, and P. G. Steeneken, Tuning the q-factor of nanomechanical string resonators by torsion support design, Applied Physics Letters 122, 013501 (2023).
- Hoch et al. [2022] D. Hoch, X. Yao, and M. Poot, Geometric tuning of stress in predisplaced silicon nitride resonators, Nano Letters 22, 4013 (2022).
- Høj et al. [2022] D. Høj, U. B. Hoff, and U. L. Andersen, Ultra-coherent nanomechanical resonators based on density phononic crystal engineering, arXiv preprint arXiv:2207.06703 (2022).
- Sadeghi [2021] P. Sadeghi, Study of high-Q nanomechanical silicon nitride resonators, Ph.D. thesis, Wien (2021).
@@
- Gisler et al. [2022] T. Gisler, M. Helal, D. Sabonis, U. Grob, M. Héritier, C. L. Degen, A. H. Ghadimi, and A. Eichler, Soft-clamped silicon nitride string resonators at millikelvin temperatures, Physical Review Letters 129, 104301 (2022).
- Høj et al. [2021] D. Høj, F. Wang, W. Gao, U. B. Hoff, O. Sigmund, and U. L. Andersen, Ultra-coherent nanomechanical resonators based on inverse design, Nature communications 12, 5766 (2021).
- Turvey [1990] K. Turvey, An undergraduate experiment on the vibration of a cantilever and its application to the determination of young’s modulus, American Journal of Physics 58, 483 (1990).
- Bückle et al. [2021] M. Bückle, Y. S. Klaß, F. B. Nägele, R. Braive, and E. M. Weig, Universal length dependence of tensile stress in nanomechanical string resonators, Physical Review Applied 15, 034063 (2021).
- Wilson et al. [2009] D. J. Wilson, C. A. Regal, S. B. Papp, and H. Kimble, Cavity optomechanics with stoichiometric sin films, Physical review letters 103, 207204 (2009).
- Allen et al. [1999] S. M. Allen, E. L. Thomas, and R. A. Jones, The structure of materials, Vol. 44 (Wiley New York, 1999).
- Kazmerski [2012] L. Kazmerski, Polycrystalline and amorphous thin films and devices (Elsevier, 2012).
- Hansen [2004] N. Hansen, Hall–petch relation and boundary strengthening, Scripta materialia 51, 801 (2004).
- Wu et al. [2022] S. Wu, Z. Kou, Q. Lai, S. Lan, S. S. Katnagallu, H. Hahn, S. Taheriniya, G. Wilde, H. Gleiter, and T. Feng, Dislocation exhaustion and ultra-hardening of nanograined metals by phase transformation at grain boundaries, nature communications 13, 5468 (2022).
- Gottstein and Shvindlerman [2009] G. Gottstein and L. S. Shvindlerman, Grain boundary migration in metals: thermodynamics, kinetics, applications (CRC press, 2009).
- Ritchie [2011] R. O. Ritchie, The conflicts between strength and toughness, Nature materials 10, 817 (2011).
- Kurotani and Tanaka [2022] Y. Kurotani and H. Tanaka, Fatigue fracture mechanism of amorphous materials from a density-based coarse-grained model, Communications Materials 3, 67 (2022).
- Gludovatz et al. [2013] B. Gludovatz, M. D. Demetriou, M. Floyd, A. Hohenwarter, W. L. Johnson, and R. O. Ritchie, Enhanced fatigue endurance of metallic glasses through a staircase-like fracture mechanism, Proceedings of the National Academy of Sciences 110, 18419 (2013).
- Ghadimi et al. [2017] A. H. Ghadimi, D. J. Wilson, and T. J. Kippenberg, Radiation and internal loss engineering of high-stress silicon nitride nanobeams, Nano letters 17, 3501 (2017).
- Schmid et al. [2016] S. Schmid, L. G. Villanueva, and M. L. Roukes, Fundamentals of nanomechanical resonators, Vol. 49 (Springer, 2016).
- Hernandez et al. [2014] R. R. Hernandez, S. Easter, M. L. Murphy-Mariscal, F. T. Maestre, M. Tavassoli, E. B. Allen, C. W. Barrows, J. Belnap, R. Ochoa-Hueso, S. Ravi, et al., Environmental impacts of utility-scale solar energy, Renewable and sustainable energy reviews 29, 766 (2014).
- Macho-Stadler et al. [2015] E. Macho-Stadler, M. Elejalde-García, and R. Llanos-Vázquez, Oscillations of end loaded cantilever beams, European Journal of Physics 36, 055007 (2015).
@@
- Nakao et al. [2008] S. Nakao, T. Ando, L. Chen, M. Mehregany, and K. Sato, Mechanical characterization of sic film at high temperatures by tensile test, in 2008 IEEE 21st International Conference on Micro Electro Mechanical Systems (IEEE, 2008) pp. 447–450.
@@
where , and is the real component of Young’s modulus, Poisson ratio and thickness of the substrate (silicon wafer in our case), respectively, and is the thickness of the a-SiC thin film. Apart from film stress, Young’s modulus , Poisson ratio and density of a-SiC are most relevant among all material properties to designing a-SiC resonator with targeted resonant frequency and stress distribution, which can be measured by patterning and then suspending the thin film as squared membrane, cantilevers and strings of different lengths . After suspended, the nanomechanical resonators are measured with Laser Doppler Vibrometer (LDV, Polytec PSV-400), while they are placed in a vacuum chamber pumped down to 10 mbar vacuum environment. After measuring the resonant frequencies of the squared membranes with lengths varying from 200 to 2000 um, we fit the measured data with the analytical formula for fundamental mode [105]
@@
Supporting Information (H): Fabrication process
Low pressure chemical vapor deposition (LPCVD) non-stoichiometric a-SiC films were used in this paper, deposited with different gas flow ratios (GFR) between SiHCl and 5% CH in H (GFR=2,3,4), at various deposition pressures (170/600 mTorr), and on both silicon and fused silica substrates (Table 1). The variation in deposition parameters allows us to systematically characterize the mechanical properties of LPCVD a-SiC. All a-SiC films were deposited at a temperature 760C for the same period of time (3 hours 47 minutes) to avoid film property differences caused by thermal effects and to ensure that the SiC films were composited of amorphous form instead of poly-crystalline form [73, 118].
After LPCVD a-SiC deposition, the wafers are diced into smaller chips. The chips are then exposed to electron beam lithograph to create desired patterns on the e-beam resist coated on top. Subsequently, these patterns are transferred into the a-SiC films using CHF anisotropic plasma etching. Next, the patterned chips are cleaned with dimethylformamide and Piranha solution, followed by the undercut of the silicon substrate or fused silica substrate using cryogenic SF isotropic plasma etching or vapor hydrofluoric acid. Finally, the designed a-SiC nanomechanical resonators are fabricated.
Supporting Information (I): Ringdown Measurement with Homodyne Detection

### v2

@@ Acknowledgements.
- [52] L. Barboni, G. Gillich, C. Chioncel, C. Hamat, I. Mituletu, In IOP Conference Series: Materials Science and Engineering, volume 416. IOP Publishing, 2018 012063.
- [53] V. A. Chirikov, D. M. Dimitrov, Y. S. Boyadjiev, Procedia Manufacturing 2020, 46 87.
- [54] S. Chen, Ultrasonics 2000, 38, 1-8 206.
- [55] L. G. Villanueva, S. Schmid, Physical review letters 2014, 113, 22 227201.
- [56] S. Wang, Z. Shan, H. Huang, Advanced Science 2017, 4, 4 1600332.
- [57] A. Banerjee, D. Bernoulli, H. Zhang, M.-F. Yuen, J. Liu, J. Dong, F. Ding, J. Lu, M. Dao, W. Zhang, et al., Science 2018, 360, 6386 300.
@@
- [58] A. Shafikov, B. Schurink, R. W. van de Kruijs, J. Benschop, W. Van den Beld, Z. S. Houweling, F. Bijkerk, Sensors and Actuators A: Physical 2021, 317 112456.
@@
- [59] M. D. Demetriou, M. E. Launey, G. Garrett, J. P. Schramm, D. C. Hofmann, W. L. Johnson, R. O. Ritchie, Nature materials 2011, 10, 2 123.
- [60] S. Zhang, Z. Li, K. Luo, J. He, Y. Gao, A. V. Soldatov, V. Benavides, K. Shi, A. Nie, B. Zhang, et al., National Science Review 2022, 9, 1 nwab140.
- [61] T. Namazu, IEEJ Transactions on Electrical and Electronic Engineering 2023, 18, 3 308.
- [62] J. Grunenberg, Angewandte Chemie International Edition 2001, 40, 21 4027.
- [63] L. G. Villanueva, S. Schmid, Physical review letters 2014, 113, 22 227201.
@@
- [64] S. A. Fedorov, N. J. Engelsen, A. H. Ghadimi, M. J. Bereyhi, R. Schilling, D. J. Wilson, T. J. Kippenberg, Physical Review B 2019, 99, 5 054107.
- [65] T. Bagci, A. Simonsen, S. Schmid, L. G. Villanueva, E. Zeuthen, J. Appel, J. M. Taylor, A. Sørensen, K. Usami, A. Schliesser, et al., Nature 2014, 507, 7490 81.
- [66] J. D. Teufel, T. Donner, M. Castellanos-Beltran, J. W. Harlow, K. W. Lehnert, Nature nanotechnology 2009, 4, 12 820.
- [67] J. Guo, R. Norte, S. Gröblacher, Physical review letters 2019, 123, 22 223602.
- [68] G. I. González, P. R. Saulson, The Journal of the Acoustical Society of America 1994, 96, 1 207.
- [69] Y. Tsaturyan, A. Barg, E. S. Polzik, A. Schliesser, Nature nanotechnology 2017, 12, 8 776.
@@
- [70] Z. Li, M. Xu, R. A. Norte, A. M. Aragón, F. Van Keulen, F. Alijani, P. G. Steeneken, Applied Physics Letters 2023, 122, 1 013501.
- [71] D. Høj, U. B. Hoff, U. L. Andersen, arXiv preprint arXiv:2207.06703 2022.
- [72] P. Sadeghi, Ph.D. thesis, Wien, 2021.
@@
- [73] T. Gisler, M. Helal, D. Sabonis, U. Grob, M. Héritier, C. L. Degen, A. H. Ghadimi, A. Eichler, Physical Review Letters 2022, 129, 10 104301.
@@
- [74] S. Nakao, T. Ando, L. Chen, M. Mehregany, K. Sato, In 2008 IEEE 21st International Conference on Micro Electro Mechanical Systems. IEEE, 2008 447–450.
- [75] D. Shin, A. Cupertino, M. H. de Jong, P. G. Steeneken, M. A. Bessa, R. A. Norte, Advanced Materials 2022, 34, 3 2106248.
- [76] D. J. Wilson, C. A. Regal, S. B. Papp, H. Kimble, Physical review letters 2009, 103, 20 207204.
- [77] M. Bückle, Y. S. Klaß, F. B. Nägele, R. Braive, E. M. Weig, Physical Review Applied 2021, 15, 3 034063.
- [78] E. Macho-Stadler, M. Elejalde-García, R. Llanos-Vázquez, European Journal of Physics 2015, 36, 5 055007.
- [79] P. G. Steeneken, R. J. Dolleman, D. Davidovikj, F. Alijani, H. S. Van der Zant, 2D Materials 2021, 8, 4 042001.
- [80] A. Castellanos-Gomez, R. van Leeuwen, M. Buscema, H. S. van der Zant, G. A. Steele, W. J. Venstra, Advanced Materials 2013, 25, 46 6719.
- [81] F. Duvigneau, S. Koch, R. Orszulik, E. Woschke, U. Gabbert, Tech Mech 2016, 36, 3 180.
- [82] D. Høj, F. Wang, W. Gao, U. B. Hoff, O. Sigmund, U. L. Andersen, Nature communications 2021, 12, 1 5766.
- [83] J. Pereira, L. E. Pichon, R. Dussart, C. Cardinaud, C. Y. Duluard, E.-H. Oubensaid, P. Lefaucheux, M. Boufnichel, P. Ranson, Applied Physics Letters 2009, 94, 7.
- [84] A. Kaur, P. Chahal, T. Hogan, IEEE Electron Device Letters 2015, 37, 2 142.
- [85] J. McKenna, J. Patel, S. Mitra, N. Soin, V. Švrček, P. Maguire, D. Mariotti, The European Physical Journal-Applied Physics 2011, 56, 2 24020.
- [86] X. Yu, B. Ding, H. Lu, Y. Huo, Q. Peng, X. Xiu, C. Zhang, C. Yang, S. Jiang, B. Man, et al., Journal of Alloys and Compounds 2019, 794 518.
- [87] P. P. Dey, A. Khare, SN Applied Sciences 2020, 2 1.
SUPPLEMENTARY INFORMATION
@@
where , and is the real component of Young’s modulus, Poisson ratio and thickness of the substrate (silicon wafer in our case), respectively, and is the thickness of the a-SiC thin film. Apart from film stress, Young’s modulus , Poisson ratio and density of a-SiC are most relevant among all material properties to designing a-SiC resonator with targeted resonant frequency and stress distribution, which can be measured by patterning and then suspending the thin film as squared membrane, cantilevers and strings of different lengths . After suspended, the nanomechanical resonators are measured with Laser Doppler Vibrometer (LDV, Polytec PSV-400), while they are placed in a vacuum chamber pumped down to 10-7 mbar vacuum environment. After measuring the resonant frequencies of the squared membranes with lengths varying from 200 to 2000 um, we fit the measured data with the analytical formula for fundamental mode [76]
@@
A2. Aspect ratio requirement for membrane resonators
Adapting from the analysis shown for plates under tension in the papers [79, 80], the total resonant frequency of squared membrane can be approximated as
here is the fundamental frequency of the square membrane-like resonator [76],
and is the fundamental frequency of the square plate-like resonator with clamped boundary condition is [81]
where is the length, is the thickness, is the tensile stress, is the density, is the Young’s modulus, and is the Poisson ratio of the square resonator.
To answer the reviewer’s comment: Equation (S2) is valid with high accuracy, if the requirement
is met. To make it make clear, this put a requirement on the aspect ratio of the membrane resonators:
In our case, using properties of a-SiC170 thin film as an example, the membranes fabricated with it should have dimensions that satisfy , which is true even with the smallest membranes we used in this work, . With the above argument, we show the use of equation (S2) is valid for our structures.
A3. Methodology verification with LPCVD Si3N4 with various thicknesses
To show the strength of the approach, we have redone the entire characterization for 5 different thicknesses to show that we can get results in line with what is known about a-SiN. We specifically deposited LPCVD stoichiometric silicon nitride to make sure the material composition was consistent between films, including the same furnace and deposition parameters. It should be noted that a-SiN thin-films’ properties can vary between different furnaces, deposition recipes, testing methodologies, etc. To our knowledge, this is one of the more comprehensive measurements of Young’s modulus, density, and Poisson ratio done for LPCVD stoichiometric Si3N4. These values tend to agree with values used in other rigorous studies of Si3N4 [64] which uses values of GPa, kg/m3, , GPa. We thank the reviewer for this suggestion as we feel it anchors the approach to a well-known material.
| Name | Substrate | (nm) | (MPa) | (kg/m3) | (GPa) | |
| Si3N4-1 | Silicon | 50.04 | 1060 | 3159 | 0.255 | 259 |
| Si3N4-2 | Silicon | 68.38 | 1080 | 3097 | 0.245 | 259 |
| Si3N4-3 | Silicon | 89.56 | 1080 | 3072 | 0.232 | 261 |
| Si3N4-4 | Silicon | 210.77 | 1080 | 3024 | 0.214 | 265 |
| Si3N4-5 | Silicon | 339.30 | 1080 | 2981 | 0.218 | 264 |
A4. Thickness uniformity and mechanical property consistency of a-SiC film across the wafer
To verify the film thickness and stress measurement on wafer-level is appropriate for the a-SiC thin films we studied in this work, we have conducted a number of additional experiments. Ellipsometry measurements were performed across the wafer to verify the uniformity of the a-SiC thin films’ thickness. Additionally, we assessed the mechanical properties of the a-SiC thin film using dies from various parts of the wafer as shown in Figure S6.
Figure S6 shows the uniformity in thickness across the wafer, substantiating the robustness of our measurement approach. The depicted color bar and measurement spots provide a detailed view of the a-SiCR2 thin film thickness at multiple points, illustrating a consistent thickness with a minor deviation of approximately 5nm over 100 mm.
Table S3 outlines the mechanical properties of a-SiCR2 thin films, tested on three distinct chips located at varied positions on the wafer. The observed minimal variations in thickness, stress, density, Young’s modulus, and Poisson’s ratio across the different chips underscore the reliability and consistency of our wafer-scale assessments.
We hope these additional experiments and data addresses the valid concerns regarding the representativeness and validity of our measurements.
| Location | Recipe | Substrate | (nm) | (MPa) | (kg/m3) | (GPa) | |
| Chip 1 | a-SiCR2 | Silicon | 78.47 | 760 | 2866 | 0.212 | 224 |
| Chip 2 | a-SiCR2 | Silicon | 78.73 | 760 | 2857 | 0.212 | 225 |
| Chip 3 | a-SiCR2 | Silicon | 79.04 | 760 | 2846 | 0.210 | 225 |
A5. Alternative method to measure mechanical properties of a-SiCR2 on fused silica substrate
In order to check the mechanical properties of a-SiCR2 thin film deposited on fused silica substrate, which is transparent thus limits the validity of the method we presented in this work, we developed an alternative procedure to characterize the a-SiCR2 thin film on fused silica.
To measure the stress on the film, we firstly coated 40 nm metal (niobium titanium nitride) on top of the a-SiCR2 film on the front side of the double polished fused silica wafer to make the surface detectable for the stress meter, then we use directional plasma etching to remove the a-SiCR2 thin film on the back side of the wafer. After measuring and recording the bending radius before and after the etching, we diced the wafer into chips, and subsequently measure the thickness of the a-SiCR2 thin film on top with the scanning electron microscopy (SEM), which is found to be found to be 821 nm. By knowing the thickness of the film, we calculate the tensile stress of the a-SiCR2 film on the back side, and infer a similar value for the film on the front side, due to the uniformity of the LPCVD deposition. This process can be seen as depositing a compressive film of the same thickness and opposite stress on both sides of the wafer. We found the average stress of the a-SiCR2 film on fused silica to be 1404 MPa. Since the etch rate of vapor HF is hard to control, we notice that there is a large undercut on the margins of the membranes (13.6 um). For the sake of accuracy, we use FEM to simulate the real shapes of the membranes, and find that with the density 2555 kg/m2, the membrane frequency can be fitted the best. The Poisson ratio of the film is then found to be 0.222 by fitting the frequency data of the strings, by applying the same method we used in the main text (Figure 2). Since cantilevers tend to collapse in our attempts to undercut with vapor HF, we resort to an alternative method to determine the Young’s modulus of the film, namely using the frequencies of the torsional modes of the 1d PnC nanostring.
We firstly use the mechanical parameters we obtained above to fit the frequencies of the flexural modes up to mode number 20th. By comparing the measurement data and the FEM simulation data, we find a very good match between the two, with only deviation found, as shown in Figure S7(c). Later on, we fit the Young’s modulus of the material by changing it slightly in the FEM simulation, which allows us to obtain the that can fit the data of torsional mode frequency optimally, in our case GPa. Note that the frequency of the torsional modes instead of the flexural modes is used, because it is more sensitive to the change of . As shown in [2], the frequency of the -th torsional mode of a uniform doubly-clamped string, with high aspect ratio , is
Here is the width of the string resonator. Unlike the frequency of the flexural mode,
the frequency of torsional mode depending more strongly on , since its second term in the bracket is not negligible, i.e. in our case. However, the torsional mode frequency of a corrugating string cannot be modelled properly with the above equation, thus we use FEM simulation instead to obtain a good fitting to the data. This alternative method allows us to measure of the material with high accuracy ( GPa). Between the simulation results obtained with GPa and the data, the deviation is less than 0.2% for all torsional modes. All in all, the alternative procedure we demonstrate above allows us to determine the relevant mechanical properties of a-SiCR2 on fused silica.
A6. Plotting the frequency dependence on device lengths with different scales
Different plots show different features in our case. Linear plots are clearer for the readers to comprehend the physical dimensions of the devices. Plots with 1/L and 1/L2 can show the higher frequency data points clearer, but its clarity about the scale of the devices might be a bit lower for the readers, they are plotted in Figure S9; Alternatively log scale plot can show clearly the length dependence of different devices, and show the physical dimensions of the devices easily. We plot them in Figure S8.
@@
We can analytically calculate the Q factor of a phononic crystal (PnC) nanostring because its losses are dominated by bending-losses from the material and does not include clamping losses to the substrate. However, simple double-clamped strings have modes which couple directly to the substrate, and thus the mechanical loss is now dependent on both the nanostring material and the substrate, which make difficult to use quality factor of simple nanostrings to intrinsic Q of the material (. As is shown in [16], for simple doubly-clamped string, other external loss mechanism will influence the accuracy on measuring . By using the 1d PnC nanostring instead, our mode of interest has no coupling with the substrate. Moreover, using phononic crystal nanostrings we can greatly enhance the dilution factor , and thus a small difference in will be amplified into a noticeable difference in the quality factor .The Q factor dependence on frequency/length, for a specific material with a constant , is equivalent to checking the dilution factor dependence on frequency/length.
@@
F2. XPS measurement on the nano-thick a-SiC films
Thank you for bringing this up. Indeed about a nanometer of the film will be native oxide, thus for the nano- a-SiC films (4.9 to 6 nm), it is necessary to show it still consist of silicon and carbon chemical components and Si-C bonds. We performed XPS measurement on three different spots: (1) 6-nm-thick a-SiCR2 film without experiencing SF6 isotropic etching, (2) 4.9-nm-thick a-SiCR2 film after SF6 etching, (3) and 4.9-nm-thick a-SiCR2 membrane (1.5x1.5 mm2) after SF6 etching. The results of the XPS measurement is shown in Figure S16. From the spectrum shown in Figure S16(a), peaks correspond to silicon, carbon and oxygen components are measured for all three spot. Fluorine peak is found for films experienced the SF6 cryogenic etching, this was also observed in [83].
In Figure S16(b) and (c), detailed XPS spectra around Si(2p) and C(1s) are shown, respectively. In Figure S16(b), the XPS spectrum of 6-nm-thick film without SF6 cryo etching is fitted (dash purple line), with Si-C bond (101.2eV, purple solid line), Si-O-C/Si-O bond (102.7eV, cyan solid line) and SiO2 bond (103.7eV, yellow solid line) [84]. The slight difference between the spectrum of the films before and after etching, may come from Si-F and/or C-F bonds. In Figure S16(c), the XPS spectrum of 6-nm-thick film without SF6 cryo etching is fitted (dash purple line), with Si-C bond (285.3eV, purple solid line), C-O bond (287.6, cyan solid line) and C-C bond (284.3eV, yellow solid line) [85]. From the XPS spectrum, the Si-C bond is identified in all three measurement spots, indicating that nano-thick SiC films still consist if a-SiC. This plot is also added to the SI for a better insight of the nano-thick a-SiC thin film.
@@
Supporting Information (H): Chemical composition characterization of the a-SiC films
There is a necessity to perform a detailed examination and discussion on the film composition and chemical bond structure in relation to their mechanical properties. Here we have undertaken a series of meticulous analyses, including XRD, EDX, and Raman spectroscopy, to gain deeper insights into these aspects.
In order to make the characterization results more comparable, the a-SiCR2/170/R3/R4 thin films used for EDX and Raman spectroscopy, are of thicknesses 78/86/81/137 nm, respectively.
The XRD analyses (Fig S19a) revealed no discernible peaks related to crystal or poly-crystalline structures, underscoring the amorphous nature of the films [86]. A noteworthy observation from our Raman spectroscopy (Fig S19c) is the correlation between higher strength and a lower ratio of Si-Si/C-C bonds. This finding aligns with established knowledge [43]; generally, C-C single bonds, with a bond dissociation energy of approximately 348 kJ/mol (3.61 eV), are stronger than Si-Si single bonds, which possess a bond dissociation energy of around 226 kJ/mol (2.34 eV) [48].
We have integrated these additional insights and discussions into the concluding sections of our manuscript, and believe this enriches the paper by not only providing readers with a nuanced understanding of the properties of these films but also offering intuitive guidance on reproducing and optimizing these films further. These additions to the conclusion aim to clarify the compositional attributes that confer the mechanical properties observed in our a-SiC films.
Supporting Information (I): Fabrication process
Low pressure chemical vapor deposition (LPCVD) non-stoichiometric a-SiC films were used in this paper, deposited with different gas flow ratios (GFR) between SiH2Cl2 and 5% C2H2 in H2 (GFR=2,3,4), at various deposition pressures (170/600 mTorr), and on both silicon and fused silica substrates (Table 1). The variation in deposition parameters allows us to systematically characterize the mechanical properties of LPCVD a-SiC. All a-SiC films were deposited at a temperature 760C for the same period of time (3 hours 20 minutes) to avoid film property differences caused by thermal effects and to ensure that the SiC films were composited of amorphous form instead of poly-crystalline form [43, 74].
After LPCVD a-SiC deposition, the wafers are diced into smaller chips. The chips are then exposed to electron beam lithography to create desired patterns on the e-beam resist coated on top. Subsequently, these patterns are transferred into the a-SiC films using CHF3 anisotropic plasma etching. Next, the patterned chips are cleaned with dimethylformamide and Piranha solution, followed by the undercut of the silicon substrate or fused silica substrate using cryogenic SF6 isotropic plasma etching or vapor hydrofluoric acid. Finally, the designed a-SiC nanomechanical resonators are fabricated.
Supporting Information (J): Ringdown Measurement with Homodyne Detection